Project

General

Profile

Wiki » History » Version 5

« Previous - Version 5/84 (diff) - Next » - Current version
Ken Lutz, 07/15/2016 02:23 PM


GDP Deployment Notes

BWRC Server Room

Sensors and Gateways
BBB - UHKBBB004.eecs.berkeley.edu 128.32.63.176
Rear Wall BLEE - C0:98:E5:30:00:36
Mid Wall BLEE - C0:98:E5:30:00:5D
Hot Aisle BLEE- C0:98:E5:30:00:54
BLINK - C0:98:E5:90:00:1E
Lab Rear Wall BLEE - C0:98:E5:30:00:09
Power Blade - Utility power to sump pump - C0:98:E5:70:00:8F
Power Blade - Utility power to Air Compressor - C0:98:E5:70:90
USB Microphone - mounted on BBB

Operational Notes
6-17-16 2:59PM Repaired broken Blink in server room. Was reading all zero, now detecting movement correctly.

7/11/16 Fixed powerblade on air compressor - was installed backward. Data collected after 7/11 looks correct. Prior was in error. Note that energy accumulated goes to zero when powerblade was disconnected for inspection. Good marker.

7/14/16 9:29AM Correlated spike in temperature and humidity for a short period of time across all BLEES in the server room. Probably related to someone working in the room and having propped the door open.

Cory Hall Swarm Lab and Immersion Room

Operational Notes
7/12/16 2:59PM, powerblade on Speaker camera disconnected for inspection and reconnected.
Note that accumulated energy goes to zero at this time.
7/13/16 10AM, powerblade on Immersion room primary projector re-installed, now in correct orientation. Readings from 10am on look correct.
7/14/16 9:22:58AM - Introduced an anomaly by heating the Immersion room BLEE located on the rear pillar with a hot air gun for about 4 min. Temperature and humidity impacted.
7/14/16 9:46AM - Speaker camera was left on since 7/12 2:26pm. Turned off at 7/14 9:46AM
7/15/16 - 9:40AM Primary projector powerblade seems to have a loose connection. No readings from 7/14 3:23PM until reseated at 7/15 9:40AM

Jacobs Hall

Gateways
ML-100 - swarmnuc1005.eecs.berkeley.edu 169.229.62.194
BBB - uhkbbb005.eecs.berkeley.edu 169.229.62.195

Powerblades
c0:98:e5:70:00:8a
c0:98:e5:70:00:e2
c0:98:e5:70:00:e5
c0:98:e5:70:00:d3
c0:98:e5:70:00:d7
c0:98:e5:70:00:f7
c0:98:e5:70:00:db

c0:98:e5:70:01:13
c0:98:e5:70:01:22
c0:98:e5:70:01:23
c0:98:e5:70:01:25
c0:98:e5:70:01:26
c0:98:e5:70:01:27
c0:98:e5:70:01:29
c0:98:e5:70:01:0c
c0:98:e5:70:01:2e
c0:98:e5:70:01:31

BLEE
c0:98:e5:30:00:2f
c0:98:e5:30:00:55
c0:98:e5:30:00:50
c0:98:e5:30:00:57
c0:98:e5:30:00:62
c0:98:e5:30:00:51

Blink
c0:98:e5:90:00:17
c0:98:e5:90:00:86
c0:98:e5:90:00:94
c0:98:e5:90:00:96
c0:98:e5:90:00:85
c0:98:e5:90:00:87
c0:98:e5:90:00:81

Operational Notes
Sensors installed, new batteries 7/15/16